# Modeling the amplitude and energy decay of a weakly damped harmonic oscillator using the energy dissipation rate and a simple trick

### v1

@@ Abstract
We demonstrate how to derive the exponential decrease of amplitude and an excellent approximation of the energy decay of a weakly damped harmonic oscillator. This is achieved using a basic understanding of the undamped harmonic oscillator and the connection between the damping force’s power and the energy dissipation rate. The trick is to add the energy dissipation rates corresponding to two specific pairs of initial conditions with the same initial energy. In this way, we obtain a first-order differential equation from which we quickly determine the time-dependent amplitude and the energies corresponding to each pair of considered initial conditions. Comparing the results of our model to the exact solutions and energies yielded an excellent agreement. The physical concepts and mathematical tools we utilize are familiar to first-year undergraduates.
@@ I Introduction
The damped harmonic oscillator is a topic regularly covered in physics textbooks for the first year of undergraduate studies, e.g., see [1, 2, 3]. Sometimes, only figures with a graphic representation of solutions of the corresponding equation of motion are given [1], and occasionally, analytical expressions of these solutions are given along with such figures [2, 3]. Still, none of these books show the derivation of these solutions. The reason for this is that the equation of motion of the damped harmonic oscillator is a second-order differential equation, the solution of which requires mathematical tools that students at that level are not yet familiar with. Thus, solutions are just given, and usually, only the exponential approximation of the energy decay in the weak damping limit is discussed [2, 3].
@@
On the other hand, in these same books [1, 2, 3], before getting to the part with the damped harmonic oscillator, the simple harmonic oscillator (i.e., undamped harmonic oscillator) is treated in great detail. Students know the solutions of an undamped harmonic oscillator and how to relate the amplitude and phase of the undamped solution to the initial conditions. Another important concept that students learn even before learning about undamped harmonic motion is the concept of instantaneous power and its connection to the rate of change of the system’s energy [2, 3].
@@
In this paper, we show how to model the solutions and complete behavior of the energy of a weakly damped harmonic oscillator, using only prior knowledge about the undamped harmonic oscillator and the connection between the power of the damping force and the energy dissipation rate. Of course, we must also have some physical insight into the behavior of the system we are modeling. Therefore, it would be best to start by providing students with figures of the results of carefully conducted experiments on systems described by the damped harmonic oscillator model, such as experiments with mechanical oscillatory systems [4, 5] or with an oscillating RLC circuits [6], in a weakly damped regime. Another option we choose in this paper is to show graphically the exact solutions of the equation of motion of a weakly damped harmonic oscillator as a substitute for the experimental results to gain insight into its behavior and the approximations we can use.
@@
This paper is organized into five sections. In section II, we present the theory needed for modeling a weakly damped harmonic oscillator, and we state the problem we are dealing with in this paper. In section III, we explain the approximations we use in our model and derive the exponential decrease of the amplitude and the corresponding expressions for the energies for the two pairs of initial conditions we consider. In section IV, we compare the results of our model with exact solutions and energies. In section V, we comment on a simpler version of our model and derivation, which are more suitable for high school students.
@@ II Basic theory and the statement of the problem
As an example of a damped harmonic oscillator, we consider a block of mass $m$ that oscillates under the influence of the restoring force of an ideal spring of stiffness $k$ and the damping force $F_{d}(t)=-bv(t)$, where $b$ is the damping constant and $v(t)$ is the velocity of the block. The equation of motion of this system is [2]
@@
where $x(t)$ is the block’s displacement from the equilibrium position, and $a(t)$ is its acceleration. Since $v(t)=dx(t)/dt$ and $a(t)=d^{2}x(t)/dt^{2}$, (1) is a second-order linear differential equation. The methods for finding its solutions are not given in physics textbooks for the first year of undergraduate studies, e.g., see [1, 2, 3]. For any $b\geq 0$, the energy (potential plus kinetic) of the block-spring system is given by
@@
If we put $b=0$ in (1) we get the equation of the undamped harmonic oscillator [2], with general solution $x(t)=A_{0}\cos(\omega_{0}t+\phi_{0})$, where $\omega_{0}=\sqrt{k/m}$ is the angular frequency of the undamped system, while $A_{0}$ and $\phi_{0}$ are constants (amplitude and phase) that are determined from the initial conditions. The undamped system oscillates with conserved, i.e., constant, energy [2]. For the damped systems, i.e., for $b>0$, the energy is not conserved due to the power of the damping force $P_{d}(t)=F_{d}(t)v(t)$, and the energy dissipation rate is given by [3]
@@ III Modeling of the amplitude decrease and the energy decay of weakly damped harmonic oscillator
It is clear from Fig. 1(a) and (b) that the amplitudes of $\tilde{x}_{1}(t)$ and $\tilde{x}_{2}(t)$ change very slowly in time in comparison to the oscillating parts. Therefore, the rate of change of function $f(t)$ over time has to be much slower than the rate of change of $\cos(\omega_{0}t)$ and $\sin(\omega_{0}t)$, which is of the order $\omega_{0}$. We can conclude that $|\dot{f}(t)|\ll\omega_{0}$ holds for weak damping for all time, i.e., for $t\geq 0$ (we have put absolute value because we can expect the negative derivative of $f(t)$ since the amplitude decreases with time).
@@
and, by setting $v_{1}(0)=0$, one easily gets that $\phi=\arctan(\dot{f}(0)/\omega_{0}))$ resolves this issue, but, due to the condition $|\dot{f}(0)|\ll\omega_{0}$, we can see that $\phi\ll 1$ holds. This leads us to the conclusion that $\tilde{x}_{1}(t)$ has a phase, but which is so small that it is unobservable in Fig. 1(a). Thus, adding this phase into our model would be an unnecessary complication, i.e., (4) is a valid model of $\tilde{x}_{1}(t)$.
@@
Since $E_{1}(t)+E_{2}(t)=2E_{0}f(t)^{2}$, and $v_{1}^{2}(t)+v_{2}^{2}(t)=\omega_{0}^{2}x_{0}^{2}f^{2}(t)$ (in adding the squared velocities we again neglected the terms of the order $(\dot{f}(t)/\omega_{0})^{2}$) we get the differential equation
@@
Using substitutions $F(t)=f^{2}(t)$ and $\alpha=-b/m$, (14) becomes $dF(t)/dt=\alpha F(t)$, and first year undergraduates know that the function $F(t)=e^{\alpha t}$ satisfies this differential equation [3]. Thus, we obtain
@@
as the condition of weak damping. Since $\omega_{0}=\sqrt{k/m}$, (18) can be rewritten as $b\ll 2\sqrt{mk}$, which is a common form of the weak damping condition in physics textbooks [3].
@@ IV Comparison of our model with exact results
The exact solution of differential equation (1), for general initial conditions, is given, e.g., in [6]. For a pair of initial conditions, which we considered in Fig. 1(a) and (b), the exact solutions are
@@
where we rounded numbers to four decimals, thus, even without showing a graphical comparison, it is clear that our modeled solution (4) and (5), with $f(t)=e^{-0.02\omega_{0}t}$, approximate excellently the exact solutions (21) and (22). To further test the validity, but also to gain insight into the limits of our model, in Fig. 2 we show the modeled solutions (4) and (5) for $\frac{b}{2m}=0.1\omega_{0}$, and the corresponding exact solutions (19) and (20). We see only a slight mismatch between the modeled and the exact solutions. Thus, our model describes well the exact solutions (at least) for $0<\frac{b}{2m}\leq 0.1\omega_{0}$.
@@
Figure 2: Modeled solutions $x_{1}(t)$ (red curve) and $x_{2}(t)$ (blue curve) given by (4) and (5) for $\frac{b}{2m}=0.1\omega_{0}$. The corresponding exact solutions $\tilde{x}_{1}(t)$ and $\tilde{x}_{2}(t)$ given by (19) and (20) are shown as dotted curves.
@@
The exact expression for the energy of the damped harmonic oscillator, for general initial conditions, is given, e.g., in [7]. For a pair of initial conditions, which we considered in Fig. 1(a) and (b), the exact energies are
$\tilde{E}_{1,2}(t)=E_{0}e^{-\frac{b}{m}t}\left[\cos^{2}(\omega t)\pm\frac{b}{2m\omega}\sin(2\omega t)+\frac{\omega_{0}^{2}+\left(\frac{b}{2m}\right)^{2}}{\omega^{2}}\sin^{2}(\omega t)\right]\,,$ (23)
where the upper sign in $\pm$ corresponds to $\tilde{E}_{1}(t)$, and lower sign to $\tilde{E}_{2}(t)$. In Fig. 3 we show the modeled energies and the exact energies for $\frac{b}{2m}=\{0.02\omega_{0},0.05\omega_{0},0.1\omega_{0}\}$. We can see excellent agreement between our modeled energies and the exact energies.
@@
Figure 3: (a) Modeled energy $E_{1}(t)$, i.e. (16). (b) Modeled energy $E_{2}(t)$, i.e. (17). Both figures show the behavior for $\frac{b}{2m}=\{0.02\omega_{0},0.05\omega_{0},0.1\omega_{0}\}$. The corresponding exact energies $\tilde{E}_{1}(t)$ and $\tilde{E}_{2}(t)$, given by (23), are shown as dotted curves on both figures.
V A simpler version of our model

### v2

@@ Abstract
We demonstrate how to derive the exponential decrease of amplitude and an excellent approximation of the energy decay of a weakly damped harmonic oscillator without solving the associated equation of motion and without insight into the analytical form of its solution. This is achieved using a basic understanding of the undamped harmonic oscillator and the connection between the damping force’s power and the energy dissipation rate. The trick is adding the energy dissipation rates corresponding to two specific pairs of initial conditions with the same energy. In this way, we obtain a first-order differential equation from which we get the time-dependent amplitude and the energies corresponding to each pair of considered initial conditions. Comparing the results of our model to the exact solutions and energies yielded an excellent agreement. The physical concepts and mathematical techniques we employ are well-known to first-year undergraduates.
@@ I Introduction
The damped harmonic oscillator, with damping force linear in velocity, is regularly covered in physics textbooks for the first year of undergraduate studies, e.g., see Cutnell8 ; Resnick10 ; Young2020university . Sometimes, only figures with a graphic representation of solutions of the corresponding equation of motion are given Cutnell8 , and occasionally analytical expressions of these solutions are given along with such figures Resnick10 ; Young2020university . Still, none of these books show the derivation of these solutions. The reason for this is that the equation of motion for the damped harmonic oscillator is a second-order differential equation, and finding its solution involves mathematical methods that students at that level have not yet learned. Thus, solutions are just given, and usually, only the exponential approximation of the energy decay in the weak damping limit is discussed Resnick10 ; Young2020university . Furthermore, even in physics textbooks that give a more thorough analysis of vibrations and waves at a more advanced undergraduate level, an analytical form of the damped harmonic oscillator solution is given without derivation and without explanation why the amplitude is assumed to decay exponentially with time, e.g., see Waves .
@@
On the other hand, in these same textbooks Cutnell8 ; Resnick10 ; Young2020university , before getting to the part with the damped harmonic oscillator, the simple harmonic oscillator (i.e., undamped harmonic oscillator) is treated in great detail. Students know the solutions of an undamped harmonic oscillator and how to relate the amplitude and phase of the undamped solution to the initial conditions. Another important concept that students learn even before learning about undamped harmonic motion is the concept of instantaneous power and its connection to the rate of change of the system’s energy Resnick10 ; Young2020university .
@@
In this paper, we show how to model the solutions and the energy of a weakly damped harmonic oscillator, using only prior knowledge about the undamped harmonic oscillator and the connection between the power of the damping force and the energy dissipation rate. Of course, we must also have some physical insight into the behavior of the system we are modeling. Therefore, it would be best to start by providing students with figures of the results of carefully conducted experiments on systems described by the damped harmonic oscillator model, such as experiments with mechanical oscillatory systems Wang ; Hinrichse2019 ; Gonzalez2006 or with an oscillating RLC circuits Lelas2023 , in a weakly damped regime. Another option we choose in this paper is to show graphically the exact solutions of the equation of motion of a weakly damped harmonic oscillator as a substitute for the experimental results to gain insight into its behavior and the approximations we can use.
Exponential amplitude decay and purely exponential approximation of the energy decay of weakly damped harmonic oscillator have been derived using the approximation that the amplitude remains constant over time intervals of one period in length and using the energy dissipation rate averaged over these time intervals Wang ; Kontomaris2024 . Our approach allows the amplitude to vary within the period, i.e., there is no need for time averaging, and we obtain the full behaviour of the energy decay.
@@
This paper is organized into six sections. In section II, we present the theory needed for modeling a weakly damped harmonic oscillator, and we state the problem we are dealing with in this paper. In section III, we explain the approximations we use in our model and derive the exponential decrease of the amplitude and the corresponding expressions for the energies for the two pairs of initial conditions we consider. In section IV, we compare the results of our model with exact solutions and energies. In section V, we comment on a simpler version of our model and derivation, which are more suitable for high school students. In section VI, we comment on the benefits of our approach and give suggestions on how to integrate it into teaching. In section VII, we briefly comment on how to experimentally obtain the data for figures needed in our approach.
@@ II Basic theory and the statement of the problem
As an example of a damped harmonic oscillator, we consider a block of mass $m$ that oscillates under the influence of the restoring force of an ideal spring of stiffness $k$ and the damping force $F_{d}(t)=-bv(t)$, where $b$ is the damping constant and $v(t)$ is the velocity of the block. The equation of motion of this system is Resnick10
@@
where $x(t)$ is the block’s displacement from the equilibrium position, and $a(t)$ is its acceleration. Since $v(t)=dx(t)/dt$ and $a(t)=d^{2}x(t)/dt^{2}$, (1) is a second-order linear differential equation. The methods for finding its solutions are not given in physics textbooks for the first year of undergraduate studies, e.g., see Cutnell8 ; Resnick10 ; Young2020university . For any $b\geq 0$, the energy (potential plus kinetic) of the block-spring system is given by
@@
If we put $b=0$ in (1) we get the equation of the undamped harmonic oscillator Resnick10 , with general solution $x(t)=A_{0}\cos(\omega_{0}t+\phi_{0})$, where $\omega_{0}=\sqrt{k/m}$ is the angular frequency of the undamped system, while $A_{0}$ and $\phi_{0}$ are constants (amplitude and phase) that are determined from the initial conditions. The undamped system oscillates with conserved, i.e., constant, energy Resnick10 . For the damped systems, i.e., for $b>0$, the energy is not conserved due to the power of the damping force $P_{d}(t)=F_{d}(t)v(t)$, and the energy dissipation rate is given by Young2020university
@@
Equation (3) tells us that the energy of the damped system decreases monotonically with time, i.e. $\frac{dE(t)}{dt}\leq 0$ $\forall t$, and $\frac{dE(t)}{dt}=0$ holds at the turning points, i.e. at instants when $v(t)=0$.
@@ III Modeling of the amplitude decrease and the energy decay of weakly damped harmonic oscillator
It is clear from Fig. 1(a) and (b) that the amplitudes of $\tilde{x}_{1}(t)$ and $\tilde{x}_{2}(t)$ change very slowly over time in comparison to the oscillating parts. Therefore, the rate of change of function $f(t)$ over time has to be much slower than the rate of change of $\cos(\omega_{0}t)$ and $\sin(\omega_{0}t)$, which is of the order $\omega_{0}$. We can conclude that $|\dot{f}(t)|\ll\omega_{0}$ holds for weak damping for all time, i.e., for $t\geq 0$ (we have put absolute value because we can expect the negative derivative of $f(t)$ since the amplitude decreases with time).
@@
and, by setting $v_{1}(0)=0$, one easily gets that $\phi=\arctan(\dot{f}(0)/\omega_{0}))$ resolves this issue, but, due to the condition $|\dot{f}(0)|\ll\omega_{0}$, we can see that $|\phi|\ll 1$ holds. This leads us to the conclusion that $\tilde{x}_{1}(t)$ has a phase, but which is so tiny that it is unobservable in Fig. 1(a). Thus, adding this phase into our model would be redundant, i.e., (4) is a valid model of $\tilde{x}_{1}(t)$.
@@
Since $E_{1}(t)+E_{2}(t)=2E_{0}f^{2}(t)$, and $v_{1}^{2}(t)+v_{2}^{2}(t)=\omega_{0}^{2}x_{0}^{2}f^{2}(t)$ (in adding the squared velocities we again neglected the terms of the order $(\dot{f}(t)/\omega_{0})^{2}$) we get the differential equation
@@
Using substitutions $F(t)=f^{2}(t)$ and $\alpha=-b/m$, (14) becomes $dF(t)/dt=\alpha F(t)$, and first year undergraduates know that the function $F(t)=e^{\alpha t}$ satisfies this differential equation Young2020university . Thus, we obtain
@@
as the condition of weak damping. Since $\omega_{0}=\sqrt{k/m}$, (18) can be rewritten as $b\ll 2\sqrt{km}$, which is a common form of the weak damping condition in physics textbooks Young2020university .
We note here that our trick (13) of adding the dissipation rates works only for damping force linear in velocity, i.e. for $F_{d}(t)=-bv(t)$. For other types of damping, such as damping due to sliding friction with $F_{d}\propto-v(t)/|v(t)|$ Zonetti , or damping due to air resistance with $F_{d}\propto-v(t)^{3}/|v(t)|$ Mungan , our trick does not lead to expressions that can be simplified using trigonometric identities. For all three mentioned types of damping, approximate solutions, in the limit of weak damping, can be obtained by averaging the energy dissipation rate over a period $T_{0}$ Wang , but this method gives only exponential approximation of energy decay in case $F_{d}(t)=-bv(t)$ and is mathematically more demanding than our method.
@@ IV Comparison of our model with exact results
The exact solution of differential equation (1), for general initial conditions, is given, e.g., in Lelas2023 . For a pair of initial conditions, which we considered in Fig. 1(a) and (b), the exact solutions are
@@
where we rounded numbers to four decimals, thus, even without showing a graphical comparison, it is clear that our modeled solution (4) and (5), with $f(t)=e^{-0.02\omega_{0}t}$, approximate excellently the exact solutions (21) and (22). Of course, the agreement between our modeled and exact solutions will be good only if condition (18) holds because only then $\omega=\omega_{0}\sqrt{1-\left(\frac{b}{2m\omega_{0}}\right)^{2}}\approx\omega_{0}$ is valid and the initial phase in (19) is negligible. To quantify the validity of our model, but also to gain insight into its limitations, in Fig. 2(a), we show $\omega$ as a function of ratio $\frac{b}{2m}$ and we can see that it starts to significantly deviate from $\omega_{0}$ at about $\frac{b}{2m}\approx 0.1\omega_{0}$.
Figure 2: (a) Damped angular frequency $\omega$ as a function of ratio $\frac{b}{2m}$. (b) Initial phase of solution (19) as a function of ratio $\frac{b}{2m}$.
In Fig. 2(b), we show the initial phase of the exact solution (19) as a function of $\frac{b}{2m}$ and we can see that it increases in magnitude more rapidly, with the increase of $\frac{b}{2m}$, than $\omega$ decreases. For $\frac{b}{2m}=0.1\omega_{0}$, we get $\omega=0.995\omega_{0}$ and $\arctan\left(\frac{b}{2m\omega}\right)=0.1$, i.e. the first maximum of cosine in exact solution (19) is shifted forward in time by $\Delta t=0.1/\omega=0.016T_{0}$, in comparison to the first maximum of cosine in our modeled solution (4) at $t=0$. We take these differences in angular frequency and initial phase for $\frac{b}{2m}=0.1\omega_{0}$ as the limits of acceptable deviation of the modeled solutions from the exact solutions. In that sense, we can safely say that our model describes well the exact solutions for $0<\frac{b}{2m}\lesssim 0.1\omega_{0}$.
To back up our choice visually, in Fig. 3(a) and (b) we show the modeled solutions (4) and (5) for $\frac{b}{2m}=\{0.1\omega_{0},0.3\omega_{0}\}$, and the corresponding exact solutions (19) and (20). We see only a slight mismatch between the modeled and the exact solutions for $\frac{b}{2m}=0.1\omega_{0}$, while, as expected, the discrepancy is more pronounced for $\frac{b}{2m}=0.3\omega_{0}$. Of course, due to the differences in angular frequency, the difference in phase of cosine functions in modeled solutions and exact solutions increases linearly with time, i.e. $\Delta\phi\sim(\omega_{0}-\omega)t$, but the oscillations, for $0<\frac{b}{2m}\lesssim 0.1\omega_{0}$, become negligible much sooner than this phase difference significantly increases, since the amplitude decays exponentially with time.
@@
Figure 3: Modeled solutions $x_{1}(t)$ (red curve) and $x_{2}(t)$ (blue curve) given by (4) and (5), for (a) $\frac{b}{2m}=0.1\omega_{0}$ and (b) $\frac{b}{2m}=0.3\omega_{0}$. The corresponding exact solutions $\tilde{x}_{1}(t)$ and $\tilde{x}_{2}(t)$ given by (19) and (20) are shown as dotted curves.
@@
The exact expression for the energy of the damped harmonic oscillator, for general initial conditions, is given, e.g., in Lelas2024 . For a pair of initial conditions, which we considered in Fig. 1(a) and (b), the exact energies are
$\tilde{E}_{1,2}(t)=E_{0}e^{-\frac{b}{m}t}\left[1\pm\frac{b}{2m\omega}\sin(2\omega t)+2\left(\frac{b}{2m\omega}\right)^{2}\sin^{2}(\omega t)\right]\,,$ (23)
where the upper sign in $\pm$ corresponds to $\tilde{E}_{1}(t)$, and lower sign to $\tilde{E}_{2}(t)$. Again, it is easy to see that the exact energies reduce to our modeled energies (16) and (17) in case approximations $\omega\approx\omega_{0}$ and $\left(\frac{b}{2m\omega_{0}}\right)^{2}\approx 0$ are valid. In Fig. 4 we show the modeled energies and the exact energies for $\frac{b}{2m}=\{0.02\omega_{0},0.05\omega_{0},0.1\omega_{0},0.5\omega_{0}\}$. We can see excellent agreement between our modeled energies and the exact energies for $\frac{b}{2m}=\{0.02\omega_{0},0.05\omega_{0},0.1\omega_{0}\}$, with small quantitative differences visible in Fig. 4(b) for case $\frac{b}{2m}=0.1\omega_{0}$, but even in that case the modeled energy seems to show qualitatively good behavior, i.e., it decreases monotonically with plateaus in vicinity of the turning points, as it should according to energy dissipation rate (3). In both Fig. 4(a) and (b), we can see that the modeled energies for $\frac{b}{2m}=0.5\omega_{0}$ show quantitatively and qualitatively wrong behavior, namely local maxima are visible in vicinity of the first turning point in both cases.
@@
Figure 4: (a) Modeled energy $E_{1}(t)$, i.e. (16). (b) Modeled energy $E_{2}(t)$, i.e. (17). Both figures show the behavior for $\frac{b}{2m}=\{0.02\omega_{0},0.05\omega_{0},0.1\omega_{0},0.5\omega_{0}\}$. The corresponding exact energies $\tilde{E}_{1}(t)$ and $\tilde{E}_{2}(t)$, given by (23), are shown as dotted curves on both figures.
One can ask the question, do local maxima in modeled energies (16) and (17) appear for ratios $\frac{b}{2m\omega_{0}}$ greater than some specific value or are present for all values of that ratio? The first derivatives of modeled energies (16) and (17) are
$\frac{dE_{1,2}(t)}{dt}=-\left(\frac{b}{m}\right)E_{0}e^{-\frac{b}{m}t}\left[1\pm\sqrt{1+\left(\frac{b}{2m\omega_{0}}\right)^{2}}\sin\left(2\omega_{0}t-\arctan\left(\frac{2m\omega_{0}}{b}\right)\right)\right]\,.$ (24)
In deriving (24) we used known trigonometric identity $a\sin(2\omega_{0}t)+b\cos(2\omega_{0}t)=\sqrt{a^{2}+b^{2}}\sin(2\omega_{0}t+\arctan(b/a))$. Since the amplitude of the sine function in (24) is greater than one for any $\frac{b}{2m}>0$, we can easily understand that (24) can be positive near the turning points of modeled solutions (4) and (5), i.e. modeled energies (16) and (17) have local maxima in the vicinity of the turning points, but, if the condition (18) holds, approximation $\left(\frac{b}{2m\omega_{0}}\right)^{2}\approx 0$ is valid, and in that case these local maxima are unobservable.
Figure 5: Parametric plot of modeled velocity $v_{1}(t)$, i.e. (6), vs. modeled displacement $x_{1}(t)$, i.e. (4), with time $t\in[0,4T_{0}]$ as parameter, shown as a green curve for $\frac{b}{2m}=0.1\omega_{0}$ in (a), and as a red curve for $\frac{b}{2m}=0.3\omega_{0}$ in (b). In both figures, the black circles are the corresponding parametric plots of the undamped case, and the black dotted curves are the parametric plots of exact velocities vs. exact displacements for the same initial conditions and the same damping. See text for details.
Yet another interesting visualization of the relationship between our modeled solutions and the exact solutions is provided by considering velocity vs. displacement parametric plots (with time as a parameter), which are known as phase space diagrams and are regularly used in classical mechanics ClassicalM . In Fig. 5 we show a parametric plot of modeled velocity $v_{1}(t)$ plotted as a function of modeled displacement $x_{1}(t)$ for $\frac{b}{2m}=0.1\omega_{0}$ (green curve in Fig. 5(a)) and for $\frac{b}{2m}=0.3\omega_{0}$ (red curve in Fig. 5(b)). As indicated in Fig. 5(a), time flows clockwise, and we consider $0\leq t\leq 4T_{0}$. In Fig. 6 we show modeled velocity $v_{2}(t)$ plotted as a function of modeled displacement $x_{2}(t)$, for $\frac{b}{2m}=0.1\omega_{0}$ (green curve in Fig. 6(a)) and for $\frac{b}{2m}=0.3\omega_{0}$ (red curve in Fig. 6(b)). On both Fig. 5 and Fig. 6, all dotted black curves show the corresponding exact velocities and exact displacements, i.e. velocities $\tilde{v}_{1,2}(t)=d\tilde{x}_{1,2}(t)/dt$ and displacements $\tilde{x}_{1,2}(t)$. In contrast, black circles show the corresponding parametric plot of the undamped harmonic oscillator.
Figure 6: Parametric plot of modeled velocity $v_{2}(t)$, i.e. (7), vs. modeled displacement $x_{2}(t)$, i.e. (5), with time $t\in[0,4T_{0}]$ as parameter, shown as a green curve for $\frac{b}{2m}=0.1\omega_{0}$ in (a), and as a red curve for $\frac{b}{2m}=0.3\omega_{0}$ in (b). In both figures, the black circles are the corresponding parametric plots of the undamped case, and the black dotted curves are the parametric plots of exact velocities vs. exact displacements for the same initial conditions and the same damping. See text for details.
Due to the chosen units, the square of the radius of the black circle is equal in magnitude to the energy of the undamped system in $E_{0}$ units, i.e. equal to one, on all figures. The square of the distance from the origin $(0,0)$ to points $(\tilde{x}_{1,2}(t),\tilde{v}_{1,2}(t))$ on black dotted curves is equal in magnitude to $\tilde{E}_{1,2}(t)$ in $E_{0}$ units. The square of the distance from the origin $(0,0)$ to points $(x_{1,2}(t),v_{1,2}(t))$ on green and red curves is not equal in magnitude to our modeled energies (16) and (17) in $E_{0}$ units, but is equal to (9) and (10) which contain a term of order $\left(\frac{b}{2m\omega_{0}}\right)^{2}$. Thus, the square of the distance from the origin to points on the red and green curves is an excellent approximation of our modeled energies (16) and (17) if condition (18) holds. With all that in mind, Fig. 5 and Fig. 6 give us a simultaneous visual comparison of the modeled displacements, velocities and energies with the exact displacements, velocities and energies. In Fig. 5, as noted earlier, we see that our modeled solution $x_{1}(t)$ does not agree with the exact solution in the initial velocity, i.e. the modeled solution $x_{1}(t)$ has initial velocity $v_{1}(0)=-\frac{bx_{0}}{2m}$, which is clearly visible for $\frac{b}{2m}=0.3\omega_{0}$. In Fig. 5 and Fig. 6, another interesting feature of the damped harmonic oscillator can be easily seen: the velocity does not have maxima in magnitude when the system passes through the equilibrium position, as it has in the undamped case. This fact has already been pointed out, e.g., in Hinrichse2019 ; Lelas2023 .
V A simpler approach